OpenStudy (anonymous):

3 ( 5x - 2 ) - 6x = 3 ( 3x + 2 )

OpenStudy (anonymous):

3 ( 5x - 2 ) - 6x = 3 ( 3x + 2 ) 15x-6-6x=9x+6 9x-9x-6-6 x=0

OpenStudy (anonymous):

X=0